OpenStudy (anonymous):

use the substitution method. d + e = 2 d - e = 4

OpenStudy (anonymous):

use the 2nd eq so d=4+e now use this in the 1st eq 4+e+e=2 4+2e=2 2e=-2 e=-1 now use this value in either eq to get d=3 Answer is (3, –1).

OpenStudy (anonymous):

i think so..

OpenStudy (anonymous):

was i correct

OpenStudy (anonymous):

yeah..

OpenStudy (anonymous):

cool